# KiDS-Legacy: Consistency of cosmic shear measurements and joint cosmological constraints with external probes

### v1

@@ 5 Combination with external data
Weak lensing on its own only constrains a nearly degenerate combination of $\sigma_{8}$ and $\Omega_{\rm m}$, which is commonly described in terms of the $S_{8}$ parameter, where the width of the degeneracy is reflected in the uncertainty on $S_{8}$. Combining our KiDS-Legacy cosmic shear measurements with external data sets allows us to break the degeneracy between the two parameters. In particular, spectroscopic galaxy surveys provide complementary constraints on the expansion history of the Universe through measurements of the BAO feature over a range of redshifts. While the BAO feature resides in the quasi-linear clustering regime, RSD measurements probe the growth of structure. We adopt BAO measurements from DESI DR1 (Adame et al., 2025) as well as BAO and RSD measurements from eBOSS DR16 (Alam et al., 2021), which provide a tight constraint on $\Omega_{\rm m}$. Additionally, observations of SN Ia provide an alternative method of constraining the expansion history of the Universe. This method relies on measurements of the luminosity distances as a function of redshift, which provide an independent constraint on the matter density. Here, we employ SN Ia measurements from the Pantheon+ compilation (Scolnic et al., 2022; Brout et al., 2022).
Table 5: Consistency metrics for the combination of KiDS-Legacy with external data.
@@
 [^11] The third and fourth columns report the Bayes ratio and the tension level inferred from the suspiciousness, respectively. The next three columns present the results of the tier 2 test for $\Omega_{\rm m}$, $S_{8}$, and their combination. The final two columns shows the tension level arising from the tier 3 test in terms of the $p$-value for data vector predictions for the external data data set listed in the second column inferred from the KiDS-Legacy TPD (column 8) and vice versa (column 9).
We treat the KiDS and the external data vectors as independent and therefore compute the joint likelihood by multiplying the individual likelihoods of each experiment. Additionally, we assume independence between the BAO and SN measurements and conduct a joint analysis of KiDS and Pantheon+ data in combination with DESI Y1 BAO and eBOSS DR16, respectively. In Table 5, we quantify the consistency between KiDS-Legacy and the external data sets. As can be observed in Fig. 10, BAO, RSD, and SN Ia measurements put a tight constraint on the matter density. Additionally, BAO measurements constrain the Hubble parameter by incorporating an external calibration of the absolute BAO scale. Since the external data sets are sensitive to parameters that are mostly unconstrained by cosmic shear, we find a good agreement between KiDS-Legacy and DESI, eBOSS, and Pantheon+ in all tests. In the joint analyses, we find the marginal mode and the HPDI to be
KiDS + DESI + Pantheon+: $S_{8}$ = $=\,0.818^{+0.015}_{-0.014}$ (20)
$\sigma_{8}$ = $=\,0.803^{+0.024}_{-0.021}$
$\Omega_{\rm m}$ = $=\,0.311^{+0.011}_{-0.012}$
KiDS + eBOSS + Pantheon+: $S_{8}$ = $=\,0.819^{+0.014}_{-0.015}$
$\sigma_{8}$ = $=\,0.798^{+0.023}_{-0.022}$
$\Omega_{\rm m}$ = $=\,0.315^{+0.012}_{-0.013}\;.$
The breaking of the $\sigma_{8}$-$\Omega_{\rm m}$ degeneracy results in a reduction in the uncertainty on $\sigma_{8}$ by about 72% compared to the fiducial constraint of W25. In terms of $S_{8}$, this corresponds to a 22% uncertainty reduction. We note, however, that the preferred degeneracy direction for KiDS-Legacy in the more general $\Sigma_{8}=\sigma_{8}(\Omega_{\rm m}/0.3)^{\alpha}$ parameterisation differs from the $\alpha=0.5$ assumed in the definition of $S_{8}$, as discussed in section 5.1 in W25. Using the preferred $\alpha=0.58$, we find $\Sigma_{8}=0.819^{+0.015}_{-0.013}$ in the combined analysis of KiDS-Legacy + DESI Y1 BAO + Pantheon+, which is consistent with the results of W25.
@@
Additionally, we present a joint analysis between cosmic shear measurements of DES and KiDS. This analysis was previously conducted in DES+KiDS, who combined cosmic shear data from DES Y3 and KiDS-1000 data and García-García et al. (2024), who reanalysed DES Y3, KiDS-1000, and HSC DR1 data with a common harmonic-space pipeline. Here, we adopt the ?KiDS-excised? DES Y3 2PCF data vector of DES+KiDS and update the combined analysis with our KiDS-Legacy COSEBIs measurements. We sample the parameter space with our KiDS-Legacy pipeline within our fiducial prior space and adopt the independent Gaussian priors for incorporating the uncertainty on the shear and redshift calibration in DES (see table 1 of DES+KiDS, ). We note that the DES Y3 measurements excluded the overlap region between both surveys, which are therefore considered to be independent. Following the methodology of DES+KiDS, we adopt independent IA parameters for both surveys. We employ the fiducial NLA-$M$ model for KiDS IA in our combined analysis. Since the prior on the NLA-$M$ parameters is survey-dependent and an application of the NLA-$M$ model to DES data is beyond the scope of this work, we model DES IA with the NLA-$z$ model, which is the fiducial IA model in the Hybrid analysis pipeline of DES+KiDS. However, as shown in W25, changes in the IA modelling only have a minor impact on the cosmological constraints in KiDS-Legacy and therefore we do not expect the choice of IA model to make a significant impact on the consistency analysis between KiDS and DES. The respective posteriors in the combined parameter space of $S_{8}$ and $\Omega_{\rm m}$ for KiDS-Legacy (yellow), DES Y3 (green), and their combination (pink) are illustrated in Fig. 11 and their consistency is quantified in Table 5. Overall, we find an agreement between both surveys up to $N_{\sigma}=0.81$, which is reported by the parameter space test in $\Omega_{\rm m}$. Thus, we consider the cosmological constraints from both surveys to be consistent, which is in agreement with the earlier study conducted with KiDS-1000 data. The joint analysis of KiDS-Legacy + DES Y3 yields
$S_{8}$ $=0.818^{+0.012}_{-0.014}$ (21)
$\sigma_{8}$ $=0.838^{+0.059}_{-0.060}$
$\Omega_{\rm m}$ $=0.277^{+0.042}_{-0.028}\;.$
The consistency analysis between KiDS-Legacy and various probes of the low-redshift Universe individually signifies a good agreement. Therefore, this enables a joint analysis with our compilation of external data sets. We note that we do not consider a joint analysis with both DESI Y1 BAO and eBOSS DR16 data given the overlapping survey footprints. Therefore, we conduct a joint analysis of KiDS-Legacy with DES Y3, DESI Y1 BAO, and Pantheon+ data, yielding
$S_{8}$ $=0.814^{+0.011}_{-0.012}$ (22)
$\sigma_{8}$ $=0.802^{+0.022}_{-0.018}$
$\Omega_{\rm m}$ $=0.307^{+0.011}_{-0.011}\;.$
@@
In light of the apparent $S_{8}$ tension that was reported in earlier cosmic shear studies, we quantify the consistency between our KiDS-Legacy cosmic shear constraints and CMB measurements from Planck. We conduct a consistency analysis, assuming independence between both surveys, and quantify the consistency in Table 5. For modelling the CMB power spectra we employ the CosmoPower emulator (Spurio Mancini et al., 2022), which was shown to reproduce the fiducial Planck parameter constraints with similar accuracy to common Boltzmann solvers. We find no evidence for a disagreement between KiDS-Legacy cosmic shear measurements and Planck CMB data with $N_{\sigma}=0.77$ inferred with the suspiciousness statistic. Moreover, we find $S_{8}$ to be in agreement at $0.61\sigma$ between both surveys. This removes the $S_{8}$ tension, which was found to be significant at the $\sim 2\sigma$ level in earlier KiDS studies. Combined analyses of probes of the early Universe, such as CMB measurements, and probes of the late Universe, such as cosmic shear, are commonly employed in studies of extended cosmological models beyond $\Lambda$CDM. A detailed study of extended cosmological models is beyond the scope of this work, and we therefore leave the combined analysis of KiDS-Legacy and Planck data for a forthcoming publication.
@@ 6 Summary and conclusions
In this analysis, we have demonstrated that the KiDS-Legacy cosmic shear data exhibits a high level of internal consistency. We find agreement between the cosmological constraints from: different tomographic redshift bins; the auto- and cross-correlation measurements; the COSEBIs, band power spectra and 2PCF statistics; the North and South regions within the KiDS footprint; red and blue galaxies. Our three tiers of consistency metrics use Bayesian evidence, measured shifts in multi-dimensional parameter space and translated posterior distributions to quantify consistency between different data splits. From our range of consistency tests it is worth highlighting two key results. In a red-blue galaxy split analysis, we confirm the results of previous studies that find strong intrinsic galaxy alignments between red early-type galaxies and no significant alignment in the blue galaxy population. This supports our decision to adopt a new colour-dependent NLA-M model in our primary KiDS-Legacy analysis. In our analysis of the cosmic shear signal measured across the six tomographic redshift bins, we find a $N_{\sigma}<1.39$ consistency. This represents a marked improvement over previous KiDS analyses where the second tomographic bin, covering a range of $0.3<z_{\rm B}\leq 0.5$, was identified as a significant outlier. We credit this improvement to advances in redshift calibration methodology and the enhanced spectroscopic data set adopted for KiDS-Legacy (see Wright et al., 2025a, for details).
In Wright et al. (2025b) we present our fiducial cosmic shear analysis, finding $S_{8}=0.815^{+0.016}_{-0.021}$. We use our three-tier analysis to demonstrate external consistency between this result and BAO constraints [^12] Similar results are found when combining KiDS-Legacy with BAO and RSD galaxy clustering data from eBOSS DR16, where $S_{8}=0.819^{+0.014}_{-0.015}$. from DESI DR1 combined with the Pantheon+ SN Ia compilation. We infer a joint constraint of $S_{8}=0.818^{+0.015}_{-0.014}$, representing a 22% reduction in uncertainty on this parameter over our fiducial result. KiDS-Legacy is also shown to be consistent with cosmic shear data from DES Y3, where a joint analysis of the two surveys finds $S_{8}=0.818^{+0.012}_{-0.014}$. Combining the KiDS, DES, DESI and Pantheon data sets we deliver a 1.4% precision measurement of $S_{8}=0.814^{+0.011}_{-0.012}$. This result is consistent with $S_{8}$ measurements of the cosmic microwave background by Planck, which is in $0.77\sigma$ agreement with KiDS-Legacy.
KiDS-Legacy cosmic shear data exhibits a high level of internal and external consistency resulting from significant improvements in the data reduction and scientific analyses of KiDS since its inception over a decade ago. Upcoming cosmology experiments will be required to pass the three tiers of stringent internal consistency analyses that we have presented here, and as such this analysis provides a useful blueprint for future studies.
@@ Acknowledgements.
BS, ZY, and CH acknowledge support from the Max Planck Society and the Alexander von Humboldt Foundation in the framework of the Max Planck-Humboldt Research Award endowed by the Federal Ministry of Education and Research. AHW is supported by the Deutsches Zentrum für Luft- und Raumfahrt (DLR), made possible by the Bundesministerium für Wirtschaft und Klimaschutz, and acknowledges funding from the German Science Foundation DFG, via the Collaborative Research Center SFB1491 ”Cosmic Interacting Matters - From Source to Signal”. PB acknowledges financial support from the Canadian Space Agency (Grant No. 23EXPROSS1) and the Waterloo Centre for Astrophysics. LL is supported by the Austrian Science Fund (FWF) [ESP 357-N]. SJ acknowledges the Ramón y Cajal Fellowship (RYC2022-036431-I) from the Spanish Ministry of Science and the Dennis Sciama Fellowship at the University of Portsmouth. HHo, SSL, and MY acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program with Grant agreement No. 101053992. LM acknowledges the financial contribution from the grant PRIN-MUR 2022 20227RNLY3 “The concordance cosmological model: stress-tests with galaxy clusters” supported by Next Generation EU and from the grant ASI n. 2024-10-HH.0 “Attività scientifiche per la missione Euclid – fase E”. MvWK acknowledges the support by the UKSA and STFC (grant no. ST/X001075/1). AD, HHi, CM, and RR are supported by an ERC Consolidator Grant (No. 770935). HHi is supported by a DFG Heisenberg grant (Hi 1495/5-1), the DFG Collaborative Research Center SFB1491, and the DLR project 50QE2305. MB, PJ, and AJW are supported by the Polish National Science Center through grant no. 2020/38/E/ST9/00395. MB is also supported by the Polish National Science Center through grant no. 2020/39/B/ST9/03494. LP acknowledges support from the DLR grant 50QE2002. TT acknowledges funding from the Swiss National Science Foundation under the Ambizione project PZ00P2_193352. YZ acknowledges the studentship from the UK Science and Technology Facilities Council (STFC). BJ acknowledges support by the ERC-selected UKRI Frontier Research Grant EP/Y03015X/1 and by STFC Consolidated Grant ST/V000780/1. KK acknowledges support from the Royal Society and Imperial College. SSL is receiving funding from the programme ?Netzwerke 2021?, an initiative of the Ministry of Culture and Science of the State of Northrhine Westphalia. CG is funded by the MICINN project PID2022-141079NB-C32. CH acknowledges support from the UK Science and Technology Facilities Council (STFC) under grant ST/V000594/1. MA is supported by the UK Science and Technology Facilities Council (STFC) under grant number ST/Y002652/1 and the Royal Society under grant numbers RGSR2222268 and ICAR1231094. CM acknowledges support from the Beecroft Trust, the Spanish Ministry of Science under the grant number PID2021-128338NB-I00. JHD acknowledges support from an STFC Ernest Rutherford Fellowship (project reference ST/S004858/1). MR acknowledges financial support from the INAF grant 2022”. AL acknowledges support from the research project grant ‘Understanding the Dynamic Universe’ funded by the Knut and Alice Wallenberg Foundation under Dnr KAW 2018.0067. NRN acknowledges financial support from the National Science Foundation of China, Research Fund for Excellent International Scholars (grant n. 12150710511), and from the research grant from China Manned Space Project n. CMS-CSST-2021-A01.
@@
Appendix A Catalogue level splits: data properties and B-mode analysis
In this Appendix, we summarise the data properties of the KiDS-Legacy catalogue split into mutually exclusive subsets, which are analysed in Sect. 4.2. Table 6 lists the redshift range, the fraction of sources with respect to the total number of sources, the effective number density, the ellipticity dispersion, the shift in the mean of the redshift distribution, and the multiplicative shear bias. We provide these data for the fiducial catalogue, a catalogue split by hemisphere, and two catalogues split into samples of red and blue galaxies. The redshift distribution per tomographic bin for each catalogue level split is illustrated in Fig. 12.

### v2

@@ 5 Combination with external data
Weak lensing on its own only is only equipped to constrain a nearly degenerate combination of $\sigma_{8}$ and $\Omega_{\rm m}$. This is commonly described in terms of the $S_{8}$ parameter, where the width of the degeneracy is reflected in the uncertainty on $S_{8}$. Combining our KiDS-Legacy cosmic shear measurements with external datasets allows us to break the degeneracy between the two parameters. In particular, spectroscopic galaxy surveys provide complementary constraints on the expansion history of the Universe through measurements of the BAO feature over a range of redshifts. While the BAO feature resides in the quasi-linear clustering regime, RSD measurements probe the growth of structure. We adopted BAO measurements from DESI DR1 (Adame et al., 2025) as well as BAO and RSD measurements from eBOSS DR16 (Alam et al., 2021), which provide a tight constraint on $\Omega_{\rm m}$. Additionally, observations of SN Ia provide an alternative method of constraining the expansion history of the Universe. This method relies on measurements of the luminosity distances as a function of redshift, which provide an independent constraint on the matter density. Here, we employed SN Ia measurements from the Pantheon+ compilation (Scolnic et al., 2022; Brout et al., 2022).
Table 5: Consistency metrics for the combination of KiDS-Legacy with external data. [^11] The third and fourth columns report the Bayes ratio and the tension level inferred from the suspiciousness, respectively. The next three columns present the results of the tier 2 test for $\Omega_{\rm m}$, $S_{8}$, and their combination. The final two columns shows the tension level arising from the tier 3 test in terms of the $p$-value for data vector predictions for the external data dataset listed in the second column inferred from the KiDS-Legacy TPD (column 8) and vice versa (column 9).
@@
We treated the KiDS data and the external data vectors as independent. Therefore, we computed the joint likelihood by multiplying the individual likelihoods of each experiment. Additionally, we assume independence between the BAO and SN Ia measurements and conduct a joint analysis of KiDS and Pantheon+ data in combination with DESI Y1 BAO and eBOSS DR16, respectively. In Table 5, we quantify the consistency between KiDS-Legacy and the external datasets. As can be observed in Fig. 10, BAO, RSD, and SN Ia measurements put a tight constraint on the matter density. Additionally, BAO measurements constrain the Hubble parameter by incorporating an external calibration of the absolute BAO scale. Since the external datasets are sensitive to parameters that are mostly unconstrained by cosmic shear, we found a good agreement between KiDS-Legacy and DESI, eBOSS, and Pantheon+ in all tests. In the joint analyses, we found the marginal mode and the highest posterior density interval to be
KiDS + DESI + Pantheon+: $S_{8}$ = $=0818^{+0.015}_{-0.014,}$ (20)
$\sigma_{8}$ = $=0803^{+0.024}_{-0.021,}$
$\Omega_{\rm m}$ = $=0311^{+0.011}_{-0.012,}$
KiDS + eBOSS + Pantheon+: $S_{8}$ = $=0819^{+0.014}_{-0.015,}$
$\sigma_{8}$ = $=0798^{+0.023}_{-0.022,}$
$\Omega_{\rm m}$ = $=0315^{+0.012}_{-0.013}\;.$
The breaking of the $\sigma_{8}$-$\Omega_{\rm m}$ degeneracy results in a reduction in the uncertainty on $\sigma_{8}$ by about 72% compared to the fiducial constraint of W25. In terms of $S_{8}$, this corresponds to a 22% uncertainty reduction. We note, however, that the preferred degeneracy direction for KiDS-Legacy in the more general $\Sigma_{8}=\sigma_{8}(\Omega_{\rm m}/0.3)^{\alpha}$ parameterisation differs from the $\alpha=0.5$ assumed in the definition of $S_{8}$, as discussed in sect. 5.1 in W25. Using the preferred $\alpha=0.58$, we find $\Sigma_{8}=0.819^{+0.015}_{-0.013}$ in the combined analysis of KiDS-Legacy + DESI Y1 BAO + Pantheon+, which is consistent with the results of W25.
@@
Additionally, we present a joint analysis between cosmic shear measurements of DES and KiDS. This analysis was previously conducted in DES+KiDS, who combined cosmic shear data from DES Y3 and KiDS-1000 data and García-García et al. (2024), who reanalysed DES Y3, KiDS-1000, and HSC DR1 data with a common harmonic-space pipeline. Here, we adopted the ?KiDS-excised? DES Y3 2PCF data vector of DES+KiDS and update the combined analysis with our KiDS-Legacy COSEBIs measurements. We sampled the parameter space with our KiDS-Legacy pipeline within our fiducial prior space and adopted the independent Gaussian priors for incorporating the uncertainty on the shear and redshift calibration in DES (see Table 1 of DES+KiDS, ). We note that the DES Y3 measurements excluded the overlap region between both surveys, which are therefore considered to be independent. Following the methodology of DES+KiDS, we adopted independent IA parameters for both surveys. We employed the fiducial NLA-$M$ model for KiDS IA in our combined analysis. Since the prior on the NLA-$M$ parameters is survey-dependent and an application of the NLA-$M$ model to DES data is beyond the scope of this work, we modelled DES IA with the NLA-$z$ model, which is the fiducial IA model in the Hybrid analysis pipeline of DES+KiDS. However, as shown in W25, changes in the IA modelling only have a minor impact on the cosmological constraints in KiDS-Legacy. Therefore, we did not expect the choice of IA model to make a significant impact on the consistency analysis between KiDS and DES. The respective posteriors in the combined parameter space of $S_{8}$ and $\Omega_{\rm m}$ for KiDS-Legacy (yellow), DES Y3 (green), and their combination (pink) are illustrated in Fig. 11 and their consistency is quantified in Table 5. Overall, we found an agreement between both surveys up to $N_{\sigma}=0.81$, which is reported by the parameter space test in $\Omega_{\rm m}$. Thus, we considered the cosmological constraints from both surveys to be consistent, which is in agreement with the earlier study conducted with KiDS-1000 data. The joint analysis of KiDS-Legacy + DES Y3 yields
[…]
@@
The consistency analysis between KiDS-Legacy and various probes of the low-redshift Universe individually signifies a good agreement. Therefore, this enables a joint analysis with our compilation of external datasets. We note that we did not consider a joint analysis with both DESI Y1 BAO and eBOSS DR16 data given the overlapping survey footprints. Therefore, we conduct a joint analysis of KiDS-Legacy with DES Y3, DESI Y1 BAO, and Pantheon+ data, yielding
[…]
@@
[…]
@@ 6 Summary and conclusions
In this analysis, we have demonstrated that the KiDS-Legacy cosmic shear data exhibits a high level of internal consistency. We find an agreement between the cosmological constraints from different tomographic redshift bins, the auto- and cross-correlation measurements, the COSEBIs, band power spectra and 2PCF statistics, the North and South regions within the KiDS footprint, and the red and blue galaxies. Our three tiers of consistency metrics use Bayesian evidence, measured shifts in multi-dimensional parameter space and TPDs to quantify consistency between different data splits. From our range of consistency tests, it is worth highlighting two key results. In a red-blue galaxy split analysis, we confirm the results of previous studies that find strong intrinsic galaxy alignments between red early-type galaxies and no significant alignment in the blue galaxy population. This supports our decision to adopt a new colour-dependent NLA-M model in our primary KiDS-Legacy analysis. In our analysis of the cosmic shear signal measured across the six tomographic redshift bins, we find a $N_{\sigma}<1.39$ consistency. This represents a marked improvement over previous KiDS analyses where the second tomographic bin, covering a range of $0.3<z_{\rm B}\leq 0.5$, was identified as a significant outlier. We credit this improvement to advances in redshift calibration methodology and the enhanced spectroscopic dataset adopted for KiDS-Legacy (see Wright et al., 2025a, for details).
In Wright et al. (2025b), we present our fiducial cosmic shear analysis, reporting a result of $S_{8}=0.815^{+0.016}_{-0.021}$. We used our three-tier analysis to demonstrate external consistency between this result and BAO constraints [^12] Similar results are found when combining KiDS-Legacy with BAO and RSD galaxy clustering data from eBOSS DR16, where $S_{8}=0.819^{+0.014}_{-0.015}$. from DESI DR1 combined with the Pantheon+ SN Ia compilation. We inferred a joint constraint of $S_{8}=0.818^{+0.015}_{-0.014}$, representing a 22% reduction in uncertainty on this parameter over our fiducial result. KiDS-Legacy is also shown to be consistent with cosmic shear data from DES Y3, where a joint analysis of the two surveys finds $S_{8}=0.818^{+0.012}_{-0.014}$. Combining the KiDS, DES, DESI, and Pantheon datasets, we were able to deliver a 1.4% precision measurement of $S_{8}=0.814^{+0.011}_{-0.012}$. This result is consistent with $S_{8}$ measurements of the cosmic microwave background by Planck, which is in $0.77\sigma$ agreement with KiDS-Legacy.
The KiDS-Legacy cosmic shear data exhibits a high level of internal and external consistency resulting from significant improvements in the data reduction and scientific analyses of KiDS since its inception over a decade ago. Upcoming cosmology experiments will be required to pass the three tiers of stringent internal consistency analyses presented here and, as such, this analysis provides a useful blueprint for future studies.
@@ Acknowledgements.
We thank the anonymous referee for their constructive comments, which helped to improve the manuscript. BS, ZY, and CH acknowledge support from the Max Planck Society and the Alexander von Humboldt Foundation in the framework of the Max Planck-Humboldt Research Award endowed by the Federal Ministry of Education and Research. AHW is supported by the Deutsches Zentrum für Luft- und Raumfahrt (DLR), made possible by the Bundesministerium für Wirtschaft und Klimaschutz, and acknowledges funding from the German Science Foundation DFG, via the Collaborative Research Center SFB1491 ”Cosmic Interacting Matters - From Source to Signal”. PB acknowledges financial support from the Canadian Space Agency (Grant No. 23EXPROSS1) and the Waterloo Centre for Astrophysics. LL is supported by the Austrian Science Fund (FWF) [ESP 357-N]. SJ acknowledges the Ramón y Cajal Fellowship (RYC2022-036431-I) from the Spanish Ministry of Science and the Dennis Sciama Fellowship at the University of Portsmouth. HHo, SSL, and MY acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program with Grant agreement No. 101053992. LM acknowledges the financial contribution from the grant PRIN-MUR 2022 20227RNLY3 “The concordance cosmological model: stress-tests with galaxy clusters” supported by Next Generation EU and from the grant ASI n. 2024-10-HH.0 “Attività scientifiche per la missione Euclid – fase E”. MvWK acknowledges the support by the UKSA and STFC (grant no. ST/X001075/1). AD, HHi, CM, and RR are supported by an ERC Consolidator Grant (No. 770935). HHi is supported by a DFG Heisenberg grant (Hi 1495/5-1), the DFG Collaborative Research Center SFB1491, and the DLR project 50QE2305. MB, PJ, and AJW are supported by the Polish National Science Center through grant no. 2020/38/E/ST9/00395. MB is also supported by the Polish National Science Center through grant no. 2020/39/B/ST9/03494. LP acknowledges support from the DLR grant 50QE2002. TT acknowledges funding from the Swiss National Science Foundation under the Ambizione project PZ00P2_193352. YZ acknowledges the studentship from the UK Science and Technology Facilities Council (STFC). BJ acknowledges support by the ERC-selected UKRI Frontier Research Grant EP/Y03015X/1 and by STFC Consolidated Grant ST/V000780/1. KK acknowledges support from the Royal Society and Imperial College. SSL is receiving funding from the programme ?Netzwerke 2021?, an initiative of the Ministry of Culture and Science of the State of Northrhine Westphalia. CG is funded by the MICINN project PID2022-141079NB-C32. CH acknowledges support from the UK Science and Technology Facilities Council (STFC) under grant ST/V000594/1. MA is supported by the UK Science and Technology Facilities Council (STFC) under grant number ST/Y002652/1 and the Royal Society under grant numbers RGSR2222268 and ICAR1231094. CM acknowledges support from the Beecroft Trust, the Spanish Ministry of Science under the grant number PID2021-128338NB-I00. JHD acknowledges support from an STFC Ernest Rutherford Fellowship (project reference ST/S004858/1). MR acknowledges financial support from the INAF grant 2022”. AL acknowledges support from the research project grant ‘Understanding the Dynamic Universe’ funded by the Knut and Alice Wallenberg Foundation under Dnr KAW 2018.0067. NRN acknowledges financial support from the National Science Foundation of China, Research Fund for Excellent International Scholars (grant n. 12150710511), and from the research grant from China Manned Space Project n. CMS-CSST-2021-A01.
@@
Appendix A Catalogue-level splits: Data properties and B-mode analysis
In this appendix, we summarise the data properties of the KiDS-Legacy catalogue split into mutually exclusive subsets, which are analysed in Sect. 4.2. Table 6 lists the redshift range, the fraction of sources with respect to the total number of sources, the effective number density, the ellipticity dispersion, the shift in the mean of the redshift distribution, and the multiplicative shear bias. We provide these data for the fiducial catalogue, a catalogue split by hemisphere, and two catalogues split into samples of red and blue galaxies. The redshift distribution per tomographic bin for each catalogue-level split is illustrated in Fig. 12.